# Keyed Chaotic Masking: A Functional Privacy Framework for Neural Inference

### v2

@@ Abstract
This work introduces a lightweight framework for privacy-preserving neural network inference based on keyed chaotic masking a deterministic, user-specific obfuscation method derived from cryptographically seeded chaotic dynamical systems. The approach applies masks to input and output tensors using key-conditioned graph dynamics, enabling authenticated inference, user attribution, and soft output watermarking without modifying model architectures. While the underlying chaotic system used to generate each mask is not analytically invertible, the masking operation itself is algebraically reversible by authorized key holders, offering functional privacy without formal cryptographic guarantees. Unlike traditional encryption or secure multi-party computation, this method operates in continuous space and imposes minimal computational overhead. We describe the construction of the masking system, including graph sampling, dynamical rule selection, and chaos diagnostics. Applications include privacy-preserving inference, secure data contribution, and per-user watermarking in shared model pipelines. This framework offers a practical and modular building block for user-controlled privacy in modern AI systems.
Note: This is a preliminary preprint released to share the core ideas and invite early feedback. The proposed framework uses cryptographically derived keys (e.g., via ECDH and HKDF) to seed user-specific transformations, and inherits security from these primitives. However, the chaotic transformation itself is not a formally proven cryptographic primitive. It should be viewed as a functional privacy mechanism, offering deterministic obfuscation and user attribution. Future work will explore its robustness under adversarial threat models and its relationship to formal notions of security.
@@ 1 Introduction
Modern machine learning systems increasingly operate in settings where user privacy, data security, and attribution are critical. inference pipelines often transmit sensitive user data to hosted models. In these contexts, users must trust that their data are handled securely, and that their contributions are not exposed, misused, or misattributed. Existing approaches to privacy-preserving inference often rely on heavyweight cryptographic methods such as homomorphic encryption [1], and secure multi-party computation [2], or introduce performance trade-offs through techniques like differential privacy [3] and training-time watermarking [4]. These approaches can require significant computational overhead, retraining, or infrastructure changes, limiting their practicality in real-world inference pipelines.
To the best of my knowledge, no existing method offers a lightweight, deterministic approach to tensor-level security that supports both input privacy and contribution traceability, while remaining fully compatible with standard model architectures. Ideally, such a method would act as a cryptographic wrapper for input embeddings, applied either prior to or after inference, and would support a wide range of downstream use cases including authentication, watermarking, attribution, and data contribution auditing.
This work introduces a framework for applying keyed chaotic dynamical transformations to tensors used in machine learning pipelines, enabling privacy-preserving, traceable, and non-invertible manipulation of data representations. Each transformation is derived from a cryptographically seeded dynamical system, represented as a computation graph that operates deterministically on the input. The transformation can be configured to either preserve task-relevant structure or intentionally obscure it, depending on the application, resulting in representations that are effectively non-invertible due to chaotic divergence, watermarked, and cryptographically bound to a key. These transformations can be applied generally to tensors and are designed to integrate cleanly with existing models, without modifying architecture or requiring retraining.
@@
  • A framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning pipelines, enabling user-specific, deterministic, and lightweight obfuscation of data.
  • A construction method for cryptographically seeded graph-based dynamical systems, including deterministic graph sampling, node-level chaotic update rules, and Lyapunov-based validation.
  • A conceptual demonstration of the method’s utility in input obfuscation, output watermarking, and authenticated inference, showing how these transformations support data privacy, user attribution, and secure access.
  • An exploration of downstream use cases, including privacy-preserving inference, federated data contribution, and soft watermarking for multi-user systems—highlighting the compatibility of the approach with existing model architectures.
This framework bridges ideas from cryptography, dynamical systems, and representation learning, offering a new lens on how models and data can interact securely. Unlike traditional cryptographic methods, this approach is architecture-agnostic, low-overhead, and enables per-user control over inference contributions. The following sections outline the system design, propose future experiments on secure inference and watermark robustness, and explore directions for chaos-based ML infrastructure.
@@ 2.1 Cryptography
Modern cryptographic systems rely on mathematical transformations that are deterministic, non-invertible, and key-dependent. These properties are used to ensure data confidentiality, integrity, and authenticity across a wide range of communication protocols [5]. While most cryptographic methods operate on discrete data (e.g., byte strings), the key ideas — particularly keyed deterministic transforms and collision-resistant signatures — are relevant when designing mechanisms for secure computation in neural systems.
Two foundational building blocks in this space are symmetric-key encryption and cryptographic hash functions. Symmetric schemes such as AES and HMAC rely on a shared secret key to encrypt or authenticate data, ensuring that only entities with access to the key can reproduce or verify the transformation [6, 7]. These methods are computationally efficient and deterministic, but are designed for bitwise or block-level data and do not preserve semantic or geometric structure — making them incompatible with neural network pipelines where data often lives in high-dimensional continuous tensor spaces.
@@
The method introduced in this work draws inspiration from these primitives but is fundamentally distinct: it operates in continuous space, supports tensor-valued inputs, and is task-preserving by design. By using keyed chaotic systems to transform tensors, the method introduces cryptographically inspired properties—such as key-dependent transformation and input sensitivity—while remaining lightweight and compatible with standard model architectures.
@@ 2.2 Privacy-Preserving Inference
As machine learning systems become increasingly deployed in real-world, cloud-based, and client-server settings, privacy-preserving inference has emerged as a critical area of research. In many applications, users transmit sensitive data — such as biometrics, medical signals, personal text, or proprietary feature representations — to centralized models for inference. This has raised concerns not only about data exposure, but also about the reusability, traceability, and control of these contributions once they enter the inference pipeline.
@@
The approach introduced in this work departs from these paradigms. Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior to or following inference. These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally transformed, and task-compatible (i.e., preserving shape and structure for downstream model compatibility). While the masking operation is algebraically reversible by the key holder (e.g. via subtraction), the underlying chaotic system used to generate the mask is not analytically invertible. As a result, the overall transformation remains opaque to unauthorized parties. The chaotic nature of the transformation ensures exponential sensitivity to small changes in initial state or key, rendering inversion impractical without full system knowledge—even though the transformation is technically deterministic. This enables lightweight, key-authenticated inference, where only correctly transformed inputs yield meaningful results, and offers tensor-level control over privacy, traceability, and secure data contribution—without requiring architectural changes or incurring the computational costs of conventional cryptographic methods.
@@ 2.3 Chaotic Dynamical System
Chaotic dynamical systems are deterministic processes that exhibit sensitive dependence on initial conditions, nonlinear evolution, and long-term unpredictability despite being fully governed by known rules. They are widely studied in physics, biology, and nonlinear systems theory, and have seen limited but notable use in machine learning—primarily as components in reservoir computing.
In the context of this work, chaos is used as a security primitive rather than a model architecture. Chaos is associated with the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small changes in initial conditions. The key properties of chaotic systems that make them suitable for this purpose include:
@@
  • Sensitivity to initial conditions: ensures that different inputs (or different users) produce non-overlapping, difficult-to-invert trajectories.
  • Structural complexity: allows for transformation of input tensors in ways that preserve local information but obscure global reversibility.
This work draws inspiration from systems such as coupled map lattices, chaotic attractors, and graph-structured reservoirs, using them to implement transformations that are key-conditioned, lightweight, and difficult to approximate or reverse without full knowledge of the graph topology and dynamical rules. By applying these systems to input or output tensors, it becomes possible to encode user-specific behavior, embed cryptographic structure into data, and control the visibility or traceability of representations within inference pipelines.
3 Keyed Chaotic Transformation Construction
The goal of this method is to construct a graph-based dynamical system whose behavior is key-conditioned and exhibits characteristics of chaos—including sensitivity to initial conditions, deterministic unpredictability, and structural complexity. This system serves as a transformation mechanism for input or output tensors, enabling secure and traceable inference without modifying the model architecture. The construction process consists of sampling a graph topology, selecting a dynamical update rule, evaluating chaotic behavior (e.g., via Lyapunov exponents), and deterministically assembling the system using a cryptographic key. The following subsections outline each stage in detail.
3.1 Graph Topology Sampling
@@
The chaotic system used for transformation is implemented over a graph, where nodes represent internal state variables and edges define the pattern of dynamical interaction. The topology of this graph is sampled deterministically from a cryptographic key, which seeds a pseudorandom number generator. The goal is to produce a fixed but key-specific structure that exhibits rich dynamical behavior when coupled with nonlinear update rules.
Given a cryptographically derived key $k$, a pseudorandom byte stream is generated using a deterministic key derivation function (e.g., HKDF with SHA-256). This stream is used to seed a pseudorandom number generator for deterministic graph sampling and system assembly. When required for compatibility with legacy APIs, a 32-bit integer seed may be extracted from the first four bytes of the stream.
The resulting generator is used to deterministically sample a graph $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
In this work, topologies are sampled from families such as Erdős–Rényi graphs and Watts–Strogatz small-world graphs, though the framework generalizes to other constructions. For Erdős–Rényi graphs, edges are sampled independently:
$\mathbb{P}\big{(}(i,j)\in E\big{)}=p\quad\text{for all }i\neq j.$ (1)
for a fixed edge probability $p\in(0,1)$ and edge list $E$. Optionally, edge weights $w_{ij}\in[-a,a]$ may be sampled and scaled to control spectral properties. For small-world graphs, a Watts–Strogatz procedure is used, beginning from a $k$-regular ring lattice with each edge rewired independently with probability $\beta$.
Graphs are directed, may be weighted or unweighted, and typically disallow self-loops. The number of nodes is chosen to match the dimensionality of the input tensor segment being transformed (e.g., embedding size or output dimensionality).
Preliminary observations suggest that the structure of the graph strongly influences the resulting dynamics—including stability, memory, and sensitivity to perturbation. For example, graphs with higher connectivity (or spectral radius) may produce more stable or diffusive behavior, while sparse or modular graphs may exhibit more chaotic evolution. However, a full characterization of the relationship between graph topology and dynamical behavior is left as future work and presents a rich area for both empirical and theoretical exploration.
3.2 Dynamical Rule Selection
@@
Each node in the system evolves according to a discrete-time update rule that combines a local nonlinear map with global coupling from neighboring nodes. The resulting dynamical behavior emerges from the interaction between the node-level nonlinearity and the graph structure. After this update, an optional leaky integration step is applied to blend the new state with the previous one.
@@
Let $\mathbf{x}(t)\in\mathbb{R}^{d}$ represent the state of the system at time $t$, and let $A\in\{0,1\}^{d\times d}$ and $W\in\mathbb{R}^{d\times d}$ denote the binary adjacency matrix and edge weight matrix, respectively. The state update proceeds in two stages:
3.2.1 Model-specific update rule:
$\tilde{\mathbf{x}}(t+1)=f\left(\mathbf{x}(t),A,W;\theta\right)$
@@ 3.2.1 Model-specific update rule:
where $f$ denotes the chosen node-level update rule (e.g., logistic map, Izhikevich, linear), and $\theta$ represents optional rule-specific parameters (e.g., $r$, $\gamma$, thresholds). In most cases, $f$ incorporates both local nonlinear updates and graph-mediated coupling via the weighted adjacency matrix.
3.2.2 Leaky integration (optional):
$\mathbf{x}(t+1)=(1-\alpha)\cdot\mathbf{x}(t)+\alpha\cdot\tilde{\mathbf{x}}(t+1)$
where $\alpha\in[0,1]$ is the leak rate.
@@ 3.2.2 Leaky integration (optional):
In this work, several node-level maps are supported:
  • Logistic map: $f(x_{i})=rx_{i}(1-x_{i})$, where $r\in(3.57,4)$ controls chaotic intensity.
  • Linear map: $f(x_{i})=x_{i}$, for baseline comparison.
  • Izhikevich dynamics: A two-variable spiking neuron model used to explore biologically inspired dynamics.
@@
At runtime, the update rule is selected from a library of registered dynamics, and its parameters are sampled from known regimes that exhibit chaotic behavior. This selection is deterministically conditioned on the cryptographic key, enabling the transformation system to be both modular and reproducible across users and devices.
3.3 Chaos Diagnostics and Characterization

### v3

# Keyed Chaotic Dynamics for Privacy-Preserving Neural Inference

@@ Abstract
Neural network inference typically operates on raw input data, increasing the risk of exposure during preprocessing and inference. Moreover, neural architectures lack efficient built-in mechanisms for directly authenticating input data. This work introduces a novel encryption method for ensuring the security of neural inference. By constructing key-conditioned chaotic graph dynamical systems, we enable the encryption and decryption of real-valued tensors within the neural architecture. The proposed dynamical systems are particularly suited to encryption due to their sensitivity to initial conditions and their capacity to produce complex, key-dependent nonlinear transformations from compact rules. This work establishes a paradigm for securing neural inference and opens new avenues for research on the application of graph dynamical systems in neural network security.
@@ 1 Introduction
Modern machine learning systems increasingly operate in environments where user privacy, data security, and attribution are critical. Inference pipelines often transmit sensitive user data to hosted models; in these contexts, users must trust that their data are handled securely, and that their contributions are not exposed, misused, or misattributed. Existing approaches to privacy-preserving inference often rely on heavyweight cryptographic methods such as homomorphic encryption [1], and secure multi-party computation [2] which require significant computational overhead or infrastructure changes, limiting their practicality in real-world systems and motivating the need for lightweight, deployable alternatives.
To address this gap, we seek a practical solution for securing real-valued tensors during inference—one that supports both data privacy and attribution without disrupting existing model architectures. Ideally, such a solution would act as a cryptographic wrapper for tensors, applied either prior or post inference, without incurring significant computational overhead.
This work introduces a framework for constructing chaotic graph dynamical systems and applying them to encrypt and decrypt tensors used in machine learning pipelines, enabling privacy-preserving and traceable manipulation of data representations. Each dynamical system is constructed from a cryptographic key and is represented as a computation graph that can readily be incorporated into existing neural architectures.
@@
  • A framework for constructing chaotic graph dynamical systems from cryptographic keys.
  • A demonstration of the application of chaotic graph dynamical systems for encrypting and decrypting tensors with a focus on inference pipelines.
  • A discussion of the implications of this paradigm for future research directions in privacy preserving machine learning.
This framework bridges ideas from cryptography and dynamical systems offering a new lens on how models and data can interact securely. Unlike traditional cryptographic methods, this approach is architecture-agnostic, low-overhead, and enables per-user control over inference contributions. The following sections outline the approach, initial validation experiments and directions for future work.
@@ 2.1 Cryptography
Modern cryptographic systems rely on key-dependent, invertible transformations that enable secure and authorized access to data. These transformations, when combined with additional primitives such as cryptographic hashes and digital signatures, support confidentiality, integrity, and authenticity across a wide range of secure communication protocols [3]. While most cryptographic methods operate on discrete data (e.g., byte strings), the key ideas — particularly keyed deterministic transforms and collision-resistant signatures — are relevant when designing mechanisms for secure computation in neural systems.
Two foundational building blocks in traditional cryptography are symmetric-key encryption and cryptographic hash functions. Symmetric schemes such as AES and HMAC rely on a shared secret key to encrypt or authenticate data, ensuring that only entities with access to the key can reproduce or verify the transformation [4, 5]. These methods are computationally efficient and deterministic, but are designed for bitwise or block-level data and require quantization or serialization to operate on high-dimensional tensors. While invertible, they are not natively compatible with neural network pipelines, which often operate on continuous, real-valued representations during inference.
@@
The method introduced in this work draws inspiration from these primitives but is fundamentally distinct: it operates in continuous space and supports tensor-valued inputs. By using keys to construct graph dynamical systems that encrypt tensors, the method introduces cryptographically inspired properties, such as key-dependent transformation and input sensitivity, while remaining lightweight and compatible with standard model architectures.
@@ 2.2 Privacy-Preserving Inference
As machine learning systems become increasingly deployed in real-world, cloud-based, and client-server settings, privacy-preserving inference has emerged as a critical area of research. In many applications, users transmit sensitive data, such as biometrics, medical signals, personal text, or proprietary feature representations, to centralized models for inference. This has raised concerns not only about data exposure, but also about the reusability, traceability, and control of these contributions once they enter the inference pipeline.
@@
The approach introduced in this work departs from these paradigms. Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior or post inference. These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally transformed, and task-compatible (i.e., preserving shape and structure for downstream model compatibility). While the masking operation is algebraically reversible by the key holder (e.g. via replaying forward dynamics and performing subtraction), it requires a key as with traditional encryption methods like AES. As a result, the overall transformation remains opaque to unauthorized parties. The chaotic nature of the transformation ensures exponential sensitivity to small changes in initial state or key, rendering inversion impractical without full system knowledge—even though the transformation is technically deterministic. This enables lightweight, key-authenticated inference, where only correctly transformed inputs yield meaningful results, and offers tensor-level control over privacy, traceability, and secure data contribution—without requiring architectural changes or incurring the computational costs of conventional cryptographic methods.
@@ 2.3 Chaotic Dynamical System
Chaotic dynamical systems are processes that exhibit sensitive dependence on initial conditions, nonlinear evolution, and long-term unpredictability despite often being fully governed by known rules. They are widely studied in physics, biology, and nonlinear systems theory [9], and have seen limited but notable use in machine learning—primarily as components in reservoir computing.
In the context of this work, chaos is used as a security primitive with analogies drawn between the properties of chaotic dynamics and the properties of traditional encryption algorithms. Chaos in this work is identified by the presence of a positive Lyapunov exponent [10], indicating exponential sensitivity to small changes in initial conditions. The key properties of chaotic systems that make them suitable as a security primitive include:
@@
  • Sensitivity to initial conditions: Small changes in initial state result in vastly different trajectories.
  • Structural complexity: Chaotic transformations encrypt input tensors into seemingly unstructured forms, making them difficult to interpret or invert without full system knowledge.
In traditional encryption algorithms such as AES, security is achieved through a series of key-dependent, invertible transformations—typically involving substitution-permutation networks and multiple rounds of mixing operations. While structurally different, the chaotic transformations used in this work serve a comparable role: they are deterministic, key-conditioned, and designed to obscure the relationship between input and output tensors. Like AES, small changes to the key or input yield drastically different outputs, ensuring unpredictability and resistance to unauthorized inversion. However, unlike AES, our transformations operate in continuous space, are differentiable, and integrate seamlessly into existing machine learning pipelines—enabling lightweight, task-compatible obfuscation without requiring model modification or cryptographic decoding layers.
3 Keyed Chaotic Dynamics
Figure 1: Overview of the keyed chaotic masking and unmasking process. A cryptographic key is used to instantiate a chaotic graph dynamical system, which produces a deterministic mask matrix $S$. During encryption (top path), this mask is element-wise added to the input tensor X, producing a masked tensor $\tilde{X}=X+S$. During decryption (bottom path), the same mask is regenerated using the shared key and subtracted from $\tilde{X}$ to recover the original input: $X=\tilde{X}-S$. The central graph diagram represents the chaotic system generated from the key, and the matrices represent tensor-valued data.
3.1 Key Hashing and Seed Derivation
The chaotic graph dynamical system used in this work is deterministically instantiated from a cryptographic key through a structured key schedule and parameter sampling process. We assume the presence of a secure key exchange protocol, such as Elliptic Curve Diffie-Hellman (ECDH), to derive a shared secret key $k$. To prevent reuse of masks across different inputs or sessions, we also introduce a public nonce $n$, which is combined with the key during subkey derivation. To securely and modularly configure the chaotic system from this key, we adopt a domain-specific key schedule inspired by symmetric encryption algorithms such as AES. We start by deriving multiple subkeys from the master key k by hashing together fixed domain labels and the nonce with the key:
[…]
Each resulting subkey is then used to seed its corresponding sampling procedure: graph topology construction, chaotic map parameter selection, and initial condition generation. This separation improves modularity, ensures reproducibility, and avoids parameter entanglement between system components.
3.2 Constructing Chaotic Graph Dynamical Systems
The goal of this section is to construct a graph dynamical system whose behavior is key-conditioned and exhibits characteristics of chaos. This system serves as the foundation for a transformation mechanism that operates on input or output tensors, enabling secure and traceable inference without modifying neural network architectures.
3.2.1 Graph Sampling
@@
The chaotic system used for transformation is implemented over a graph, where nodes represent internal state variables and edges determine the coupling between node dynamics. The topology of this graph is sampled deterministically from a cryptographic key, producing a fixed but key-specific structure that supports rich dynamical behavior when combined with nonlinear update rules.
Given a cryptographically derived subkey $k_{\text{graph}}$, obtained via SHA-256 of a labeled key expansion step, we treat it as a high-entropy seed for initializing a pseudorandom number generator (PRNG). This ensures deterministic graph sampling while maintaining strong key-dependent variation. If needed for compatibility with legacy APIs, a 32-bit integer seed can be extracted from the first four bytes of $k_{\text{graph}}$. The resulting PRNG instance is then used to construct the system’s graph topology $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed. In practice we represent the graph with a single matrix $W\in\mathbb{R}^{d\times d}$ representing the coupling strengths between nodes.
In this work, adjacency matrices $A$ representing graph topologies are first deterministically sampled from classical random graph families, including Erdős–Rényi (ER) and Watts–Strogatz (WS) small-world models. For ER graphs, edges are sampled independently with fixed probability $p$, while small-world graphs are generated using a $k$-regular lattice with random rewiring governed by parameter $\beta$.
Having sampled an adjacency matrix $A$, edge weights which define coupling strengths between nodes are sampled. In the simplest case, a weight matrix $W$ is drawn from a uniform distribution over a fixed interval and subsequently the Hadamard product is taken between this matrix and the adjacency matrix as follows:
$W\leftarrow W\odot A$ (2)
As with the sampling of adjacency matrices, the current framework generalizes to arbitrary weight sampling distributions, enabling finer control over spectral properties and interaction strength. The resulting key-specific graph serves as the backbone of the chaotic system, defining how local dynamics interact and evolve over time.
3.2.2 Node Dynamics
@@
Each node in the system evolves according to a discrete-time update rule that combines a local nonlinear map with global coupling from neighboring nodes. The resulting dynamical behavior emerges from the interaction between the node-level nonlinearity and the graph structure.
@@
Let $\mathbf{x}(t)\in\mathbb{R}^{d}$ represent the state of the system at time $t$, and let $W\in\mathbb{R}^{d\times d}$ denote the sampled weight matrix representing the graph. The state update proceeds according to:
$\tilde{\mathbf{x}}(t+1)=f(\mathbf{x}(t),W;\phi)=f_{\text{local}}(\mathbf{x}(t);\phi)+f_{\text{coupled}}(\mathbf{x}(t),W;\phi)$ (3)
@@
where $f$ denotes the chosen node-level update dynamics, and $\phi$ represents dynamics-specific parameters. In most cases, $f$ incorporates both local nonlinear updates and graph-mediated coupling via the weighted matrix.
In this work, several candidates for local nonlinear updates are used, including:
Map Name Update Rule Output State Parameter
Logistic map $x_{i+1}=rx_{i}(1-x_{i})$ $x_{i}\in(0,1)$
Tent map $x_{i+1}=\begin{cases}\frac{x_{i}}{\mu}&x_{i}<\mu\\ \frac{1-x_{i}}{1-\mu}&x_{i}\geq\mu\end{cases}$ $x_{i}\in(0,1)$
Baker’s map $(x_{i+1},y_{i+1})=\begin{cases}(sx_{i},y_{i})&x_{i}<s\\ ((1-s)x_{i}+s,(1-s)y_{i}+s)&x_{i}\geq s\end{cases}$ $x_{i}\in[0,1]$
[…]
Arnold’s cat map $\begin{bmatrix}x_{i+1}\\ y_{i+1}\end{bmatrix}=\begin{bmatrix}1&1\\ 1&2\end{bmatrix}\begin{bmatrix}x_{i}\\ y_{i}\end{bmatrix}\mod 1$ $x_{i}\in[0,1]$
@@
Table 1: Summary of candidate node-level chaotic maps used in this work.
Different methods of coupling between nodes can be applied to each of the local nonlinear update rules. In this work we choose diffusive coupling between nodes which is defined as:
$\tilde{x}_{i}(t+1)=f_{\text{local}}(x_{i}(t);\phi)+\sum_{j=1}^{d}W_{ij}\left(f_{\text{local}}(x_{j}(t);\phi)-f_{\text{local}}(x_{i}(t);\phi)\right)$ (4)
@@
At runtime, the update rule is selected from a library of registered dynamics, and its parameters are sampled from ranges known to result in bounded chaotic dynamics. This selection is deterministically conditioned on the cryptographic key, enabling the transformation system to be both modular and reproducible across users and devices.
3.2.3 Chaotic Dynamics Diagnostics